Calculate the following:

1.1.1(-39)0

1.1.2 (9)5/7 rounded off to 3 significant figures

1.1.3 -5 +3 ÷ 2

1.1.4 -(-32)0

1.2 Find the value for the expression (x2)3 / x5

1
Expert's answer
2022-05-25T12:31:29-0400

Calculate the following:

1.1.1(-39)= 1

1.1.2 (9)5/7 rounded off to 3 significant figures = 4.804

1.1.3 -5 +3 ÷ 2 = -3.5

1.1.4 -(-32)0 = -1

1.2 Find the value for the expression (x2)3 / x= x